OpenStudy (anonymous):

If h(x) = 2x + 3, find the value of x if h(x) = 9. 21 3 6 9

OpenStudy (anonymous):

h(x)=9 h(x)=2x+3 9=2x+3 9-3=2x 6=2x 6/2=x 3=x